OpenStudy (anonymous):

Sketch the graph of 4x-4y=2? Does any one no how to graph this help me plz thank you!!!!!!!!!!!!!! :)

jimthompson5910 (jim_thompson5910):

One way to graph this is to find the x and y intercepts

OpenStudy (anonymous):

first you will need to put it slope-intercept form

jimthompson5910 (jim_thompson5910):

x-intercept: plug in y = 0 and solve for x 4x-4y=2 4x-4*0=2 4x - 0 = 2 4x = 2 x = 2/4 x = 0.5 so the x-intercept is (0.5, 0)

OpenStudy (anonymous):

How do I do that

jimthompson5910 (jim_thompson5910):

Do you see how to find the y-intercept?

jimthompson5910 (jim_thompson5910):

If the intercepts method confuses you, then you can isolate y and place the equation in slope intercept form. So you have to subtract 4x from both sides. Then divide both sides by -4 to isolate y. Then place the equation into y = mx+b form \[\Large 4x-4y=2\] \[\Large -4y=2-4x\] \[\Large -4y=-4x+2\] \[\Large y=\frac{-4x+2}{-4}\] \[\Large y=\frac{-4x}{-4}+\frac{2}{-4}\] \[\Large y=x-\frac{1}{2}\]